OpenStudy (elise_a18):

FAN & MEDAL? The square root of 50 plus square root of 2 Decide irrational or rational and solve.

OpenStudy (campbell_st):

well to help look at it like this \[\sqrt{50} + \sqrt{2} = \sqrt{25 \times 2} + \sqrt{2}\] this may enable you to simplify it

OpenStudy (elise_a18):

does it simplify to \[\sqrt[6]{2}\]..............? .-.

OpenStudy (campbell_st):

not quite \[\sqrt{25 \times 2} + \sqrt{2} = \sqrt{25} \times \sqrt{2} + \sqrt{2}\] thismay be a bigger help

OpenStudy (campbell_st):

so you get \[5\sqrt{2} + \sqrt{2} = ? \sqrt{2}\] collect the like terms to find ?

OpenStudy (elise_a18):

To be honest. I'm not good at all with radicals. They hit me like a bag of hammers :'/

OpenStudy (elise_a18):

\[\sqrt[7]{2}\] ?? since the indexes are added?

OpenStudy (campbell_st):

no this small 7 means the 7th root of 2 when the answer is just \[6 \sqrt{2}\]

OpenStudy (elise_a18):

ohh I see now. But how did you get the 6th root of 2 instead of 7?

OpenStudy (campbell_st):

its \[\sqrt{50} = \sqrt{25 \times 2} = \sqrt{25} \times \sqrt{2} = 5 \times \sqrt{2}\] and \[\sqrt{2} = 1 \times \sqrt{2}\] so the common factor is \[\sqrt{2}\] which means the solution is \[(5 + 1) \sqrt{2} = 6 \sqrt{2}\] its like 5x + x = 6x

OpenStudy (elise_a18):

Oh okay. Sorry about that. I have to watch some videos or something on radicals. Thanks so much for your help Campbell :D